# Adjoint log canonical foliated singularities on surfaces

### v1

@@ A.2. Special $K^{>0}$-chain
The proof follows similarly to Proposition 4.5 and is in fact simpler in this setting. ∎
A.3. Classification
Let $G_{E}$ denote the dual graph of $E$, which is connected.
A.3.1. Case I: $G_{E}$ consists of a single curve
Proposition A.3.
If $G_{E}=\{[C]\}$, then $C$ is one of:
  1. (1)
    a smooth rational curve. In this case, KC=2KC=-2.
  2. (2)
    a smooth elliptic curve. In this case, $KC=0$.
  3. (3)
    a rational curve with a single node. In this case, $KC=0$.
Proof.
This follows from Lemma A.1. ∎
A.3.2. Case II: $E$ is a $K^{=0}$-graph
Proposition A.4.
If $KC=0$ for every component $C\subset E$, then $E$ is one of:
  1. (1)
    an irreducible curve: a smooth elliptic curve or a rational curve with a single node,
  2. (2)
    a cycle of smooth rational curve.
Proof.
This follows from Lemma A.1. ∎
A.3.3. The presence of special $K^{>0}$-chain
We now assume that $E$ does not belong to Case I or Case II. Then $E$ necessarily contains at least one special $K^{>0}$-chain.
Proposition A.5.
Suppose that there exists two special $K^{>0}$-chains $F_{1}$ and $F_{2}$ with distinct first curves, such that they share a common component or satisfy $F_{1}F_{2}>0$. Then $E$ is a Hirzebruch-Jung chain.
Proof.
This follows from Proposition A.2. ∎
To separate the remaining cases, we impose the following technical assumption.
    $E$ does not belong to Case I or Case II, and any two special $K^{>0}$-chains with distinct first curves are pairwise disjoint.
Under the assumption ()(*), we can define the maximal special $K^{>0}$-chain. Let $\{\,\Theta_{1},\cdots,\Theta_{s}\,\}$ denotes the set of all maximal special $K^{>0}$-chains contained in $E$. Define
$K_{1}:=K-\sum_{i=1}^{s}M(K,\Theta_{i}).$
It is clear that $E$ is a $(K_{1})^{\geq 0}$-graph and for any component $C$ of $\Theta_{i}$, one has $K_{1}C=0$.
Lemma A.6.
Under the assumption ()(*), let $C$ be an irreducible curve contained in $E$, and assume $C$ is not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
  1. (1)
    $C$ is smooth rational curve with $d(C)=3$, meeting two special $K^{>0}$-chains $\Theta_{1},\Theta_{2}$ of type $(2,2)$. In this case, $K_{1}C=0$.
  2. (2)
    $C$ is a smooth rational curve with $d(C)=2$, disjoint from all special $K^{>0}$-chain. In this case, $K_{1}C=KC=0$.
  3. (3)
    $E=C+\Theta_{1}+\cdots+\Theta_{l}$ is a single fork centered at $C$, where $C$ is a smooth rational curve and $(\Theta_{1},\cdots,\Theta_{l})$ is of type $(m_{1},\cdots,m_{l})$, which belongs to one of
    1. (a)
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
    In the case (b) and (c), $K_{1}C=0$.
Proof.
Without loss of generality, we assume that $C$ meets transversely the chains $\Theta_{1},\cdots,\Theta_{k}$. Let $m_{j}$ denots the absolute value of the determinant of the intersection matrix of $\Theta_{j}$. Note that $m_{j}\geq 2$. Thus, we have
$K_{1}C$ $=KC-\sum_{i=1}^{k}M(K,\Theta_{i})C$
$=2p_{a}(C)-2+d(C)-\sum_{i=1}^{k}\frac{1}{m_{i}}$
$=2p_{a}(C)-2+(d-k)+\sum_{i=1}^{k}(1-\frac{1}{m_{i}}),$
where $d-k\geq 0$.
If $p_{a}(C)\geq 1$, then $K_{1}C\leq 0$ implies $p_{a}(C)=1$ and $d=k=0$, contradicting our assumption. Next we assume $p_{a}(C)=0$. Then
$K_{1}C=d-k-2+\sum_{i=1}^{k}(1-\frac{1}{m_{i}}).$
  • If $d-k\geq 2$, then $K_{1}C\leq 0$ implies $k=0$ and $d=2$. In this case, $K_{1}C=KC=0$ and $C$ is a smooth rational curve with $d(C)=2$, disjoint from all special $K^{>0}$-chain. This is case (2).
  • If dk=1d-k=1, then
    $K_{1}C=-1+\sum_{i=1}^{k}(1-\frac{1}{m_{i}})\geq-1+\frac{k}{2}$
    • If $k\geq 3$, then $K_{1}C>0$.
    • If $k=2$, then $K_{1}C\leq 0$ implies $d=3$, $k=2$ and $(m_{1},m_{2})=(2,2)$. This is case (1). In this case, $K_{1}C=0$.
    • If $k=1$, then $d=2$ and $KC=0$. In this case, $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
    • If $k=0$, then $d=1$ and KC=1KC=-1. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
  • If dk=0d-k=0, then
    $K_{1}C=-2+\sum_{i=1}^{k}(1-\frac{1}{m_{i}})\geq-2+\frac{k}{2}.$
    $K_{1}C\leq 0$ implies $k\leq 4$.
    • If $k=0$, then $d=0$, contradicting our assumption.
    • If $k=1$, then $d=1$ and KC=1KC=-1. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $k=2$, then $d=2$ and $KC=0$. In this case, $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
    • If $k=3$, then $K_{1}C\leq 0$ implies
      $\frac{1}{m_{1}}+\frac{1}{m_{2}}+\frac{1}{m_{3}}\geq 1.$
      Then $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,l)$ for $l=3,4,5,6$, or $(2,4,4)$, or $(3,3,3)$. These are case (3-a) and (3-b).
      In particular, $K_{1}C=0$ for $(m_{1},m_{2},m_{3})=(2,3,6)$, or $(2,4,4)$, or $(3,3,3)$.
    • If $k=4$, then $K_{1}C\leq 0$ implies $(m_{1},\cdots,m_{4})=(2,2,2,2)$. In this case, $K_{1}C=0$. This is case (3-c).
Proposition A.7.
Under the assumption ()(*), $E$ belongs to one of the following cases:
  1. (1)
    $E$ has two forks with the dual graph as follows:
[…]
    where each component is a smooth rational curve with self-intersection at most 2-2. In this case, $p$ is not klt.
  2. (2)
    $E=C+\Theta_{1}+\cdots+\Theta_{l}$ is a single fork centered at $C$, where $C$ is a smooth rational curve and $(\Theta_{1},\cdots,\Theta_{l})$ is of type $(m_{1},\cdots,m_{l})$, which belongs to one of
    1. (a)
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
    In the case (b) and (c), $K_{1}C=0$.
Proof.
This follows from Lemma A.6. ∎
A.3.4. Summary of the classification
Theorem A.8.
Let $p$ be a log canonical singularity of a normal surface $Y$. Let
$\pi:X\to Y$
  1. (1)
    $E$ is a Hirzebruch-Jung chain.
  2. (2)
  3. (3)
    $E$ is a cycle of smooth rational curves.
  4. (4)
    $E$ has two forks with the dual graph as follows:
[…]
  5. (5)
    1. (a)
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
Here type (2), (3), (4), (5-b) and (5-c) are not klt.
Proof.
The statement follows by combining Proposition A.3, Proposition A.4, Proposition A.5, and Proposition A.7. ∎
Appendix B Log canonical foliated singularities
In this appendix, we give a proof of the classification of log canonical foliated singularities, following an approach analogous to the program developed in Section 4, differing from that of Y. Chen.
Let $(Y,\mathcal{G})$ be a foliated pair on a normal surface $Y$, and let $p\in Y$ be a log canonical singularity of $\mathcal{G}$. Let
$\pi:(X,\mathcal{F})\longrightarrow(Y,\mathcal{G})$
be a minimal resolution of $(Y,\mathcal{G},p)$, with exceptional curves $\mathfrak{E}=\bigcup_{i\in I}E_{i}$, and set
$E:=\sum_{i\in I}E_{i}.$
Then we can write
$K:=K_{\mathcal{F}}+\sum_{i\in I}\iota(E_{i})E_{i}=\pi^{*}K_{\mathcal{G}}+\sum_{i\in I}a_{i}E_{i},$
so that $E$ is a $K^{\geq 0}$-graph.
B.1. $K$-non-positive curves in $E$
Lemma B.1.
Let $C\subset E$ be an irreducible $\mathcal{F}$-invariant curve. Then
(B.1) $KC=2p_{a}(C)-2+d(C)+{\rm Z}(\mathcal{F},C)-d_{2}(C).$
Moreover, $KC\leq 0$ if and only if $C$ is one of:
  1. (1)
    $d(C)=0$:
    1. (a)
      an irreducible elliptic Gorenstein leaf. In this case, $KC=0$,
    2. (b)
      a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=0$,
    3. (c)
      a (1)(-1)-$\mathcal{F}$-curve. In this case, KC=1KC=-1.
  2. (2)
    $d(C)=1$:
    1. (a)
      a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=1$, $d_{2}=0$. In this case, $KC=0$.
    2. (b)
      a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=1$. In this case, KC=1KC=-1,
    3. (c)
      a (2)(-2)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=1$. In this case, $KC=0$.
  3. (3)
    $d(C)=2$:
    1. (a)
      a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=d_{2}=1$. In this case, $KC=0$.
    2. (b)
      a (2)(-2)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=2$. In this case, $KC=0$.
Proof.
This follows from a direct computation. ∎
Lemma B.2.
Let $C\subset E$ be a non-$\mathcal{F}$-invariant curve. Then
(B.2) $KC={\rm tang}(\mathcal{F},C)+d_{1}(C)\geq 0,$
and $KC=0$ if and only if $C$ is smooth, with ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$.
Proof.
This follows from a direct computation. ∎
B.2. Special $K^{>0}$-chains
Proposition B.3.
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with first curve $C_{1}$. Then $F$ is an $\mathcal{F}$-chain with
$d(C_{1})=d_{2}(C_{1})=1,\quad d(C_{i})=d_{2}(C_{i})=2\,(2\leq i\leq r-1),\quad d(C_{r})=d_{2}(C_{r})\in\{1,2\},$
and
$KC_{1}=-1,\quad KC_{i}=0\text{ for }i\geq 2.$
Proof.
The proof follows similarly to Proposition 4.5 and is in fact simpler in this setting. ∎
Lemma B.4.
Any two special $K^{>0}$-chains with distinct first curves are disjoint.
Proof.
This directly follows from the separatrix theorem. ∎
B.3. Classification
Let $G_{E}$ denote the dual graph of $E$, which is connected.
B.3.1. Case I: $G_{E}$ consists of a single curve
Proposition B.5.
If $G_{E}=\{[C]\}$, then $C$ is one of:
  1. (1)
    an irreducible elliptic Gorenstein leaf. In this case, $KC=0$.
  2. (2)
    a smooth non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$. In this case, $KC=0$.
  3. (3)
    a (2)(-2)-$\mathcal{F}$-curve. In this case, $KC=0$.
  4. (4)
    a (1)(-1)-$\mathcal{F}$-curve. In this case, KC=1KC=-1.
Proof.
This follows from Lemma B.1 and Lemma B.2. ∎
B.3.2. Case II: $E$ is a $K^{=0}$-graph
Proposition B.6.
If $KC=0$ for every component $C\subset E$, then $E$ is one of:
  1. (1)
    An irreducible curve: irreducible elliptic Gorenstein leaf or smooth non-$\mathcal{F}$-invariant elliptic curve with ${\rm tang}(\mathcal{F},C)=0$.
  2. (2)
    A chain of (2)(-2)-curves.
  3. (3)
    A cycle of (2)(-2)-$\mathcal{F}$-curves.
  4. (4)
    An $\mathcal{F}$-star graph centered at a smooth non-$\mathcal{F}$-invariant curve $C$ with ${\rm tang}(\mathcal{F},C)=0$.
Proof.
This follows from Lemma B.1, Lemma B.2 and the separatrix theorem. ∎
B.3.3. Remaining cases
Let $\{\Theta_{1},\dots,\Theta_{s}\}$ be the set of all maximal special $K^{>0}$-chains in $E$, and set
$K_{1}:=K-\sum_{i=1}^{s}M(K,\Theta_{i}).$
Lemma B.7.
Let $C\subset E$ be irreducible and not in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then $C$ is one of the following:
  1. (1)
    $E=C_{1}+C+C_{2}$ is a chain where $C_{1},C_{2}$ are (1)(-1)-$\mathcal{F}$-curves, and $C_{2}$ is a bad tail. In this case, $K_{1}C=0$.
  2. (2)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve meeting two (1)(-1)-$\mathcal{F}$-curves $C_{1},C_{2}$, with $d(C_{1})=d(C_{2})=1$, $C_{1}^{2}=C_{2}^{2}=-2$ and $d(C)=d_{2}(C)=3$. In this case, $K_{1}C=0$.
  3. (3)
    $C$ is smooth non-$\mathcal{F}$-invariant, ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  4. (4)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve, $d_{1}=d_{2}=1$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  5. (5)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve, $d_{1}=0$, $d_{2}=2$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  6. (6)
    $C$ is a (1)(-1)-$\mathcal{F}$-curve, $d=d_{1}=1$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
  7. (7)
    $C$ is a (2)(-2)-$\mathcal{F}$-curve, $d=d_{2}=1$, disjoint from all special $K^{>0}$-chains. In this case, $K_{1}C=KC=0$.
Proof.
If $C$ is not $\mathcal{F}$-invariant, then it is disjoint from all special $K^{>0}$-chains. Hence $K_{1}\cdot C=K\cdot C\geq 0$, with equality if and only if $C$ is smooth, ${\rm tang}(\mathcal{F},C)=d_{1}(C)=0$, and $C$ is disjoint from all special $K^{>0}$-chains. This corresponds to case (3).
We now assume that $C$ is $\mathcal{F}$-invariant. Without loss of generality, suppose that $C$ meets $\Theta_{1},\dots,\Theta_{k}$ transversely at distinct points $p_{1},\dots,p_{k}$. Let $n_{i}$ denote the absolute value of the determinant of the intersection matrix of $\Theta_{i}$. Note that $n_{i}\geq 2$ for all $i$. Then
$K_{1}\cdot C$ $=K\cdot C-\sum_{i=1}^{k}M(K,\Theta_{i})\cdot C$
$=2p_{a}(C)-2+d(C)+{\rm Z}(\mathcal{F},C)-d_{2}(C)-\sum_{i=1}^{k}\frac{1}{n_{i}}$
$=2p_{a}(C)-2+(d-k)+({\rm Z}(\mathcal{F},C)-d_{2})+\sum_{i=1}^{k}\Bigl(1-\frac{1}{n_{i}}\Bigr),$
where $d\geq k$ and ${\rm Z}(\mathcal{F},C)\geq d_{2}$.
If $p_{a}(C)\geq 1$, then $K_{1}\cdot C\leq 0$ forces $p_{a}(C)=1$ and $d=k=0$, contradicting our assumption.
We therefore assume $p_{a}(C)=0$. In this case,
$K_{1}\cdot C$ $=-2+(d-k)+({\rm Z}(\mathcal{F},C)-d_{2})+\sum_{i=1}^{k}\Bigl(1-\frac{1}{n_{i}}\Bigr)$
$\geq-2+(d-k)+({\rm Z}(\mathcal{F},C)-d_{2})+\frac{k}{2}$
By the separatrix theorem, we have ${\rm Z}(\mathcal{F},C)\geq k+1$.
  • If $d-k\geq 3$, then $K_{1}\cdot C\geq 1>0$.
  • If dk=2d-k=2, then $K_{1}\cdot C\leq 0$ implies $k=0$, $d=2$ and ${\rm Z}(\mathcal{F},C)=d_{2}\in\{1,2\}$. Hence $C$ is either a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$, or a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=2$. In both cases, $C$ is disjoint from all special $K^{>0}$-chains and $K_{1}C=KC=0$. This yields case (4) or (5).
  • Suppose dk=1d-k=1.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}\geq 2$, then $K_{1}\cdot C>0$.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}=1$, then $K_{1}\cdot C\leq 0$ implies $k=0$, $d=1$ and ${\rm Z}(\mathcal{F},C)=d_{2}+1\in\{1,2\}$. Hence $C$ is either a (1)(-1)-$\mathcal{F}$-curve with $d=d_{1}=1$, or a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=1$. Both cases satisfy $K_{1}C=KC=0$ and are disjoint from all special $K^{>0}$-chains. This corresponds to case (6) or (7).
    • If ${\rm Z}(\mathcal{F},C)-d_{2}=0$, then
      $k+1\leq{\rm Z}(\mathcal{F},C)=d_{2}\leq d=k+1,$
      so $d=d_{2}={\rm Z}(\mathcal{F},C)=k+1$.
        If $k\geq 3$, then $K_{1}C>0$.
        If $k=2$, then $K_{1}C\leq 0$ forces $(n_{1},n_{2})=(2,2)$ and $K_{1}\cdot C=0$, giving case (2).
        If $k=1$, then $K_{1}C\leq 0$ implies that $C$ is a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=2$. In this case $KC=0$, so $\Theta_{1}+C$ forms a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
        If $k=0$, then $K_{1}C=-1$, so $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
  • If dk=0d-k=0, then $1\leq d_{2}=d=k<{\rm Z}(\mathcal{F},C)$.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}\geq 2$, then $K_{1}C>0$.
    • If ${\rm Z}(\mathcal{F},C)-d_{2}=1$, then
      $K_{1}C=-1+\sum_{i=1}^{k}\Bigl(1-\frac{1}{n_{i}}\Bigr)\geq-1+\frac{k}{2}.$
      Hence $K_{1}C\leq 0$ implies $k\in\{1,2\}$. If $k=1$, then $KC=0$ and $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$. If $k=2$, this gives case (1).
Proposition B.8.
If $E$ does not belong to Case I or II, then $E$ is one of:
  1. (1)
    An $\mathcal{F}$-chain $C_{1}+\cdots+C_{r}$.
  2. (2)
    A chain $C_{1}+C+C_{2}$ where $C_{1},C_{2}$ are (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 and $C_{2}$ is a bad tail. In this case, $p$ is not log terminal.
  3. (3)
    An $\mathcal{F}$-dihedral graph. (See Figure 5.)
Proof.
This follows directly from Lemma B.7 together with the separatrix theorem. ∎
Conclusion
We now summarize the classification.
Theorem B.9 (Y. Chen).
Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface with log canonical singularity at $p$, and let $\pi:(X,\mathcal{F})\to(Y,\mathcal{G})$ be the minimal resolution with exceptional divisors $E=\cup_{i}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 3.)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.18.)
  3. (3)
  4. (4)
    a $\mathcal{F}$-dihedral graph: two (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 joined by a bad tail connecting to a chain of (2)(-2)-$\mathcal{F}$-curves. (See Figure 5.)
  5. (5)
  6. (6)
    An $\mathcal{F}$-star chain: a chain with exactly one non-$\mathcal{F}$-invariant curve $E_{k}$, ${\rm tang}(\mathcal{F},E_{k})=0$, with the remaining curves forming $\mathcal{F}$-chains. (See Figure 8.)
  7. (7)
Type (1) is terminal; types (1)–(5) are canonical.
Proof.
Combine Propositions B.5, B.6, and B.8. ∎

### v2

@@ 5. $\epsilon$-adjoint canonical singularities for $\epsilon\in(0,1/4)$
It follows directly from Theorem 5.7, together with the classification of lc surface singularities (Theorem 2.24). ∎
5.1. Application to adjoint MMP
We now explain how the $\epsilon$-adjoint canonical classification yields a natural threshold for the adjoint minimal model program.
Let $(X,\mathcal{F})$ be a foliated surface, where $X$ is smooth and $\mathcal{F}$ is a rank one foliation with canonical singularities. Fix $\epsilon\in(0,1/4)\cap\mathbb{Q}$. We now run an $\epsilon$-adjoint MMP for $(X,\mathcal{F})$.
Choose a maximal adjoint collection of $(K_{\mathcal{F}}+\epsilon K_{X})^{\geq 0}$-graphs:
$\{\,\mathfrak{E}_{1},\cdots,\mathfrak{E}_{s}\,\}.$
Such a collection may not be unique if $K_{\mathcal{F}}+\epsilon K_{X}$ is not big.
By contracting $\mathfrak{E}_{1},\cdots,\mathfrak{E}_{s}$, we obtain a $(K_{\mathcal{F}}+\epsilon K_{X})$-non-positive birational contraction
$\varphi:(X,\mathcal{F})\longrightarrow(Y,\mathcal{G}).$
The pair $(Y,\mathcal{G})$ satisfies the following properties:
  • (a)
    For any irreducible curve $C\subset Y$, if
    $(K_{\mathcal{G}}+\epsilon K_{Y})\cdot C\leq 0,$
    then $C^{2}\geq 0$.
  • (b)
    $(Y,\mathcal{G})$ has at most $\epsilon$-adjoint canonical singularities. By Corollary 5.9, it follows that $Y$ has klt (hence rational) singularities and that $\mathcal{G}$ has lc singularities.
Consequently, we distinguish two cases:
  • (i)
    If $K_{\mathcal{F}}+\epsilon K_{X}$ is pseudo-effective, then $K_{\mathcal{G}}+\epsilon K_{Y}$ is nef. In particular, if $K_{\mathcal{F}}+\epsilon K_{X}$ is big, then $K_{\mathcal{G}}+\epsilon K_{Y}$ is ample. Since $Y$ has rational singularities, this follows from [Tan04, Theorem 4.8]. Moreover, [Tan04] provides an explicit number $\alpha=\alpha(K_{\mathcal{G}}+\epsilon K_{Y})$ such that $\alpha\cdot(K_{\mathcal{G}}+\epsilon K_{Y})$ is very ample (see also [LWX25, Corollary 1.2]).
  • (ii)
    If $K_{\mathcal{F}}+\epsilon K_{X}$ is not pseudo-effective, then neither is $K_{\mathcal{G}}+\epsilon K_{Y}$. Hence there exists an irreducible curve $C\subset Y$ with $C^{2}\geq 0$ and
    $(K_{\mathcal{G}}+\epsilon K_{Y})\cdot C<0.$
    In particular,
    $K_{Y}\cdot C<0\quad\text{or}\quad K_{\mathcal{G}}\cdot C<0.$
    By [KM98, Corollary 3.17] and [SS23, Theorem 2.8], there exists a morphism $f:Y\to Z$ such that $\rho(Y/Z)=1$ and $(K_{\mathcal{G}}+\epsilon K_{Y})$ is $f$-ample.
6. Key Examples
In this section, we present two illustrative examples.
Example 6.1.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$, and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=ny\,\mathrm{d}x-x\,\mathrm{d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over $p$ (see Figure 7).
$\cdots\cdots$
$\bar{E}_{2}(-2)$
$\bar{E}_{1}(-2)$
$\bar{E}_{n-1}(-2)$
$E_{n}(-1)$
Figure 7.
Denote by $q_{i}$ (resp. $E_{i}$), $i=1,\dots,n$, the successive blow-up points (resp. exceptional curves). In this case, we have
In particular, each $E_{i}$ is $\mathcal{F}^{\prime}$-invariant for $i\leq n-1$, while $E_{n}$ is not $\mathcal{F}^{\prime}$-invariant. A straightforward computation yields
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{n},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+\cdots+(n-1)\bar{E}_{n-1}+nE_{n}.$
Hence,
$K_{\mathcal{F}^{\prime}}+\epsilon K_{X^{\prime}}=\sigma^{*}(K_{\mathcal{F}}+\epsilon K_{X})+\epsilon\sum_{i=1}^{n-1}i\bar{E}_{i}+(n\epsilon-1)E_{n}.$
For any $\epsilon\in(0,1)$, one can choose $n$ sufficiently large so that $n\epsilon-1\geq 0$. In this situation, the point $p$ is $\epsilon$-adjoint canonical but not foliated canonical.
Remark 6.2.
Example 6.1 corresponds to type (5) in Theorem 5.7, which is the only configuration in which an $\epsilon$-adjoint canonical singularity fails to be foliated canonical for sufficiently small $\epsilon$.
Example 6.3.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$, and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=x\,\mathrm{d}x+y^{2}\,\mathrm{d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over $p$ (see Figure 8).
$\bar{E}_{1}(-3)$
$\bar{E}_{2}(-2)$
$E_{3}(-1)$
Figure 8.
Denote by $q_{i}$ (resp. $E_{i}$) the successive blow-up points (resp. exceptional curves). In this case,
$l(q_{1})=l(q_{2})=1,\qquad l(q_{3})=a(q_{3})=2.$
In particular, all $E_{i}$ are $\mathcal{F}^{\prime}$-invariant. A direct computation gives
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{3},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+4E_{3}.$
Hence,
Remark 6.4.
Example 6.3 corresponds to type (III-1-a) in Theorem 4.15, i.e., the boundary configuration that is not foliated log canonical.